Calculate the power output in watts and horsepower of a shot-- putter who takes 1.20s to accelerate the 7.27kg shot from rest to 14.0m/s, while raising it 0.800m. (5 marks)


1
Expert's answer
2021-10-18T11:14:06-0400

According to the second Newton's law, the force that acted on the shot was:


"F = ma"

where "m=7.27kg" and "a" is the acceleration:


"a = \dfrac{v_f}{t}"

where "v_f = 14m\/s" and "t = 1.20s".

The work done by the shot-putter is then:


"W = Fd = \dfrac{mv_fd}{t}"

where "d = 0.800m". By definition, the power (of a constant force) is given as follows:


"P = \dfrac{W}{t} = \dfrac{mv_fd}{t^2}\\
P = \dfrac{7.27kg\cdot 14m/s\cdot 0.8m}{(1.2s)^2} \approx 5.65\times 10^3W"

In horsepower:


"P \approx 7.69 hp"

Answer. "5.65\times 10^3W" or "7.69 hp".